A 65 ?kg human is exposed to 32.3 rad of radiation.

a)How much energy was absorbed by the person's body? Express your answer using two significant figures.

b) Compare this energy to the amount of energy absorbed by a person's body if he or she jumped from a chair to the floor (assume that the chair is 0.50 m from the ground and that all of the energy from the fall is absorbed by the person). Express your answer using two significant figures.
Answer #1

a) Mass of human = 65 kg

Radiation absorb = 32.3 rad

But 1 rad = 0.01J/ kg

\therefore 32.3 rad = 32.3* 0.01J/kg = 0.323J/kg

Energy absorbed = 0.323J / kg * 65kg = 20.995 J \approx 21J

b) E = mgh

where m= mass of person = 65kg

g = . gravity =9.8ms-2

h = 0.50m

E = 65 kg* 9.8 ms-2 * 0.50 m = 318.5 kg m2s-2

Now, 1kg m2s-2 = 1 J

E = 318.5 J